OpenStudy (civicsiscool44):

Out of the 120 students at a summer camp, 72 signed up for canoeing. There were 23 students who signed up for trekking, and 13 of those students also signed up for canoeing. Use a two-way table to organize the information and answer the following question: Approximately what percentage of students signed up for neither canoeing nor trekking? 10% 12% 38% 32%

OpenStudy (civicsiscool44):

@mathstudent55

OpenStudy (bishoppatton):

You have 120 kids in all and 23 signed up for trekking so automatically take 23 away from the 120

OpenStudy (bishoppatton):

so now there are 97

OpenStudy (civicsiscool44):

Ok

OpenStudy (dayakar):

total students = 120 no , of canoeing students = 72+13 = 85 no ,of trekking students =23

OpenStudy (civicsiscool44):

Ok your kind of confusing me

OpenStudy (dayakar):

remaining students = 120-85-23 =12

OpenStudy (civicsiscool44):

ok

OpenStudy (bishoppatton):

take away the 72 for canoeing. the 13 in the problem means nothing because it didnt ask how many slots for each activity was filled. it just asks for how many didn't do anything. so pretend you only have 26 kids trekking and 72 canoeing

OpenStudy (bishoppatton):

he is wrong

OpenStudy (bishoppatton):

dont use his answer

OpenStudy (civicsiscool44):

ok

OpenStudy (bishoppatton):

72+26=98 kids that did an activity so 22 didn't sign up for anything

OpenStudy (civicsiscool44):

ok

OpenStudy (bishoppatton):

so what percentage is 22 of 120

OpenStudy (civicsiscool44):

545?

OpenStudy (bishoppatton):

not even close

OpenStudy (dayakar):

percentage = 12/120 * 100 = 10%

OpenStudy (civicsiscool44):

is that correct????????

OpenStudy (bishoppatton):

yes